OpenStudy (anonymous):

A bag contains 12 red marbles and 4 blue marbles. What is the ratio, written in lowest terms, of red marbles to the total number of marbles in the bag? A. 1 to 4 B. 1 to 3 C. 2 to 3 D. 3 to 4

OpenStudy (anonymous):

12+4=16 12:16 12/4 = 3/4 16/4 Final answer is 3:4 I think